ON THE ENERGY DEPENDENCE OF THE ZEEMAN AND HYPERFINE PARAMETERS IN THE A^{2}Σ^{+} STATE OF OH AND OD

J. XIN, Department of Physics and Engineering Technologies, Bloomsburg University, Bloomsburg, PA 17815; I. IONESCU, D. KUFFEL, AND \underline{S. A. REID}, Department of Chemistry, Marquette University, Milwaukee, WI USA 53233.

We report quantum beat studies of the energy dependence of the Zeeman parameters in the A^{2}Σ^{+} state of OH and OD, and the nuclear hyperfine parameters of OD(A^{2}Σ^{+}). In contrast to a previous study, we find that the sign and magnitude of the anisotropic g-factor g_{l} for ν′=0 is consistent with that expected from Curl’s relationship. However, at higher energies g_{l} changes sign, and possible explanations for this trend will be examined. The magnetic and electric quadrupole hyperfine constants were determined for OD(A^{2}Σ^{+},ν′ = 0-3), and the constants for ν′ = 2,3 are reported here for the first time to our knowledge. The derived values are consistent with available ab initio calculations.